# Einstein from Noise: Statistical Analysis

### v2

@@ Abstract
“Einstein from noise” (EfN) is a prominent example of the model bias phenomenon: systematic errors in the statistical model that lead to spurious but consistent estimates. In the EfN experiment, one falsely believes that a set of observations contains noisy, shifted copies of a template signal (e.g., an Einstein image), whereas in reality, it contains only pure noise observations. To estimate the signal, the observations are first aligned with the template using cross-correlation and then averaged. Although the observations contain nothing but noise, it was recognized early on that this process produces a signal that resembles the template signal! This pitfall was at the heart of a central scientific controversy about validation techniques in structural biology.
This paper provides a comprehensive statistical analysis of the EfN phenomenon above. We show that the Fourier phases of the EfN estimator (namely, the average of the aligned noise observations) converge to the Fourier phases of the template signal, thereby explaining the observed structural similarity. Additionally, we prove that the convergence rate is inversely proportional to the number of noise observations and, in the high-dimensional regime, to the Fourier magnitudes of the template signal. Moreover, in the high-dimensional regime, the EfN estimator converges to a scaled version of the template signal. This work not only deepens the theoretical understanding of the EfN phenomenon but also highlights potential pitfalls in template matching techniques and emphasizes the need for careful interpretation of noisy observations across disciplines in engineering, statistics, physics, and biology.
@@ 1 Introduction
While the EfN phenomenon has been analyzed in prior work (see ’previous work’ in Section 2 for more details), a comprehensive theoretical understanding of the EfN model remains limited. This work contributes to filling that gap by rigorously analyzing the relationship between the reconstructed signal and the underlying template. The authors of the original article presenting the EfN phenomenon chose an image of Einstein as the template signal, and hence the name [40]. Consequently, we refer to the average of the aligned pure noise signals as the EfN estimator. The problem is formulated in detail in Section 3, and is illustrated in Figure 1.
Figure 1: Einstein from Noise. The EfN estimator consists of three stages: (1) finding the index of the maximum of the cross-correlation ($\hat{\mathsf{R}}_{i}$) between the $i$-th noise signal ($n_{i}$) and the template signal (e.g., Einstein’s image); (2) cyclically shifting the noise signal by $-\hat{\mathsf{R}}_{i}$; (3) averaging the shifted noise signals. In this paper, we characterize the relationship between the output of this process—the EfN estimator—and the template signal.
@@
The central results of this work are as follows. Our first result, stated in Theorem 4.1, shows that the Fourier phases of the EfN estimator converge to the Fourier phases of the template signal, as the number of noisy observations (denoted by $M$) converges to infinity. We also show that the corresponding mean squared error (MSE) decays to zero with a rate of $1/M$. Since the Fourier phases are responsible for the formation of geometrical image elements, such as contours and edges [33, 41], this clarifies why the resulting EfN estimator image exhibits a structural similarity to the template, but not necessarily a full recovery. Our second result, stated in Theorem 4.3, proves that in the high-dimensional regime, where the dimension of the signal diverges, the convergence rate of the Fourier phases is inversely proportional to the square of the Fourier magnitudes of the template signal. In this case, the Fourier magnitudes of the EfN estimator converge to a scaled version of the template’s Fourier magnitudes.
While Theorems 4.1 and 4.3 are proved under the assumption of white Gaussian noise, we also extend our analysis to more general noise models. In particular, we show that, although the convergence results in Theorems 4.1 and 4.3 do not necessarily hold, several structural properties of the EfN estimator persist under arbitrary noise statistics. First, in Proposition 5.1, we show that the EfN estimator remains positively correlated with the template for arbitrary noise statistics, even when the Fourier phases do not converge. Since the correlation between images often implies visual resemblance, this explains why the EfN estimator still exhibits structural similarity to the template. Second, in Theorem 5.2, we show that in the high-dimensional limit, if the noise signal is independent and identically distributed (i.i.d) (not necessarily Gaussian), then the same phase convergence behavior observed in the white Gaussian case still holds. Finally, in Proposition 5.4, we demonstrate that if the noise signal is Gaussian with circular symmetry, then the conclusions of Theorem 4.1 remain valid, even though the noise is not white.
@@
The remainder of this paper is organized as follows. The next section discusses the connection between the EfN problem and single-particle cryo-electron microscopy (cryo-EM), the primary motivation for this work, and presents supporting empirical demonstrations. Section 3 provides a detailed formulation of the problem. Our main theoretical results, Theorems 4.1 and 4.3, are stated in Section 4 and proved in Appendices C and F, respectively. Extensions of these results to noise models beyond white Gaussian noise are presented in Section 5. The key results in this section include Proposition 5.1, proved in Appendix G; Theorem 5.2, proved in Appendix H; and Proposition 5.4, proved in Appendix I. Finally, we conclude with a discussion and outlook in Section 6.
@@ 2 Cryo-EM and Empirical Demonstration
Cryo-EM is a powerful tool of modern structural biology, offering advanced methods to visualize complex biological macromolecules with ever-increasing precision. One of its central advantages lies in its capability to resolve the structures of proteins that are difficult to crystallize in traditional methods, especially in a near-physiological environment (see, e.g., [32, 44]). This advantage enables researchers to delve into the dynamic behaviors of proteins and their complexes, shedding light on fundamental biological processes.
Cryo-EM uses single-particle electron microscopy to reconstruct 3D structures from 2D tomographic projection images [7]. Typically, the 3D reconstruction involves two main steps: detecting and extracting single particle images using a particle picking algorithm, [39, 20, 9, 19], and then reconstructing the 3D density map [38, 37]. Most detection algorithms use template-matching techniques, which can introduce bias if improper templates are chosen, especially in low signal-to-noise ratio (SNR) conditions, which is the standard scenario in cryo-EM.
@@
The EfN phenomenon depends on several key parameters: (1) the number of observations denoted by $M$; (2) the dimension of the signal, denoted as $d$; and (3) the power spectral density (PSD) of the template signal. To demonstrate the dependency on these parameters and provide insight into our main results, Figures 2 and 3 show the convergence of the EfN estimator. Specifically, Figure 2 illustrates the behavior of the Fourier phases as a function of $M$. Figure 2(c) highlights that the convergence rate is proportional to $1/M$. As can be seen, the convergence rate is faster for higher spectral components. Figure 3 illustrates the impact of the PSD of the template signal on the cross-correlation between the template and the EfN estimator. Notably, a flatter PSD (i.e., a faster decay of the auto-correlation) leads to a higher correlation between the template and the estimator signals. These empirical results are proved theoretically in Theorems 4.1 and 4.3.
@@
The EfN phenomenon extends to various applications employing template matching, whether through a feature-based or direct template-based approach. For instance, template matching holds significance in computational anatomy, where it aids in discovering unknown diffeomorphism to align a template image with a target image [13]. Other areas include medical imaging processing [1], manufacturing quality control [3], and navigation systems for mobile robots [25]. This pitfall may also arise in the feature-based approach, which relies on extracting image features such as shapes, textures, and colors to match a target image by neural networks and deep-learning classifiers [54, 31, 49, 27].
@@
The EfN phenomenon has been investigated in earlier studies. In particular, it was shown that the ratios between the expected values of the Fourier coefficients of the EfN estimator and those of the template are real-valued [55, Chapter 5]. In this work, we build upon and significantly extend these results. Specifically, we establish the convergence of the EfN estimator to a non-vanishing signal, derive its convergence rate, analyze its behavior in the high-dimensional regime, and generalize the analysis to encompass a broader class of noise models beyond white Gaussian noise.
@@
A closely related work is that of Wang et al. [51], who conducted a rigorous statistical analysis of model bias in a different but complementary setting. They analyze the effects of selectively averaging only samples that exhibit the highest cross-correlation with a fixed reference signal (e.g., Einstein’s image). This selection mechanism introduces a bias toward the reference, and their analysis reveals a phase transition in the resulting reconstruction, governed by the number of samples, the signal dimension, and the size of the selected subset. Notably, their results show that a structured image can emerge even when averaging purely noisy data. In contrast, our work investigates the behavior of the EfN estimator when all the aligned pure-noise observations are averaged without any selection step. We focus on the asymptotic behavior of the Fourier phases and show that they converge to those of the underlying template signal.
Figure 2: The impact of the number of noise observations on the EfN estimator. (a) The structural similarity between the EfN estimator and the template image as a function of the number of noise observations ($M$). (b) The MSE between the phases of the template image $\mathsf{X}[k_{1},k_{2}]$ and the estimator $\hat{\mathsf{X}}[k_{1},k_{2}]$ for $-100\leq k_{1},k_{2}\leq 100$, with varying observation sizes ($M=200,500,1500,5000$). More observations lead to lower MSE. (c) The convergence rate of the MSE between the Fourier phases of the EfN estimator and the Fourier phases of the template as a function of the number of observations across different frequencies. The relationship between MSE and $M$ is $\propto M^{-1}$. In addition, stronger spectral components lead to lower MSE. Figures (b) and (c) were generated through 200 Monte-Carlo trials.
Figure 3: The influence of the PSD of the template signal on the correlation between the template and the EfN estimator. (a) Three images of the letter A are shown, with an increasing zero-padding ratio. As the zero-padding ratio increases, the PSD flattens, and the cross-correlation (CC) between the template and the EfN estimator increases. This higher cross-correlation is evident in both the image background and the colors of the letter A. (b) Flatter PSDs lead to EfN estimators whose Fourier magnitudes are closer to those of the template image. The EfN estimators in these experiments were generated using $M=10^{5}$ observations.
3 Problem Formulation and Notation
Notations.
Throughout the rest of this paper, we use ${{\cal D}}$, ${{\cal P}}$, ${\mathsf{a.s.}}$, and ${{\cal L}^{p}}$, to denote the convergence of sequences of random variables in distribution, in probability, almost surely, and in ${\cal L}^{p}$ norm, respectively. Inner products in the Euclidean space between vectors $a$ and $b$ are written as either $a^{\top}b$ or $\langle a,b\rangle$.
Problem formulation.
[…]
where $\mathcal{T}_{\ell}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the cyclic shift operator defined by $[\mathcal{T}_{\ell}z]_{r}\triangleq z_{(r-\ell)\bmod d}$ for all $z\in\mathbb{R}^{d}$ and indices $10\leq r\leq d-$ 1, and $n_{i}\sim\mathcal{N}(0,\sigma^{2}I_{d\times d})$ are i.i.d. Gaussian noise vectors. In reality, however, there is no underlying signal: the observations consist entirely of white Gaussian noise. That is, the true data-generating process follows the underlying model:
[…]
where $M$ denotes the number of observations. Without loss of generality, we assume the signal length $d$ is even. Since the data consists purely of white Gaussian noise, we will explicitly write $y_{i}=n_{i}$ to emphasize this fact.
$\hat{\mathsf{R}}_{i}\triangleq\underset{0\leq\ell<d}{\arg\max}\,\langle n_{i},\mathcal{T}_{\ell}x\rangle.$ (3.3)
$\hat{x}\triangleq\frac{1}{M}\sum_{i=0}^{M-1}\mathcal{T}_{-\hat{\mathsf{R}}_{i}}n_{i}.$ (3.4)
The EfN phenomenon refers to the empirical observation that $\hat{x}$ and $x$ often appear “close”, despite the fact that the observations contain no signal. The goal of this paper is to explain and rigorously analyze this surprising behavior.
In this work, we consider two asymptotic regimes: the first corresponds to the classical setting where the number of observations $M\to\infty$ while the dimension $d$ is fixed; the second is the high-dimensional regime, where $d\to\infty$ after $M\to\infty$.
Assumptions.
Throughout this paper, we assume that the template signal $x$ is normalized, i.e., $\left\|x\right\|_{2}^{2}=1$, where $\left\|\cdot\right\|_{2}$ is the Euclidean norm, and further assume that its Fourier transform in non-vanishing, except possibly at the DC (zero-frequency) component. The first assumption is used for convenience and does not alter (up to a normalization factor) our main results in Theorems 4.1 and 4.3. The second assumption is essential for the theoretical analysis of the EfN process and is expected to hold in many applications, including cryo-EM. A similar assumption is frequently taken in related work, e.g., [6, 35, 8]. It is worth noting that since the Fourier transform of $x$ is assumed to be non-vanishing, the maximizing shift $\hat{\mathsf{R}}_{i}$ in (3.3) is almost surely unique.
Fourier space notation.
where $j\triangleq\sqrt{-1}$, and $10\leq k\leq d-$ 1. Accordingly, we let $\mathsf{X}$, $\hat{\mathsf{X}}$, and $\mathsf{N}_{i}$, denote the DFTs of $x$, $\hat{x}$, and $n_{i}$, respectively, for $10\leq i\leq M-$ 1. These DFT sequences can be equivalently represented in the magnitude-phase domain as follows,
for $10\leq i\leq M-$ 1. Note that the random variables $\left\{|\mathsf{N}_{i}[k]|\right\}_{k=0}^{d/2}$ and $\left\{\phi_{\mathsf{N}_{i}}[k]\right\}_{k=0}^{d/2}$ are two independent sequences of i.i.d. random variables, such that, $|\mathsf{N}_{i}[k]|\sim\mathsf{Rayleigh}\left({\sigma^{2}}\right)$ has Rayleigh distribution, and the phase $\phi_{\mathsf{N}_{i}}[k]\sim\mathsf{Unif}[-\pi,\pi)$ is uniformly distributed over $[-\pi,\pi)$.
for $k=0,1,\ldots,d-1$. It is important to note that the location of the maximum correlation, i.e., $\mathsf{\hat{R}}_{i}$, captures the dependency on the template signal, as well as the connections between the different spectral components. We denote by $\mathbb{E}{\left|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]\right|^{2}}$ the MSE of the Fourier phases of the $k$-th spectral component.

### v3

@@ Abstract
“Einstein from noise” (EfN) is a prominent example of the model bias phenomenon, where systematic errors in the statistical model lead to spurious but consistent estimates. In the EfN experiment, one falsely believes that a set of observations contains noisy, shifted copies of a template signal (e.g., an Einstein image), whereas in reality, it contains only pure noise observations. To estimate the signal, the observations are first aligned with the template using cross-correlation and then averaged. Although the observations contain nothing but noise, it was recognized early on that this process produces a signal that resembles the template signal! This model bias pitfall was at the heart of a central scientific controversy about validation techniques in structural biology.
This paper provides a comprehensive statistical analysis of the EfN phenomenon above. We show that the Fourier phases of the EfN estimator (namely, the average of the aligned noise observations) converge to the Fourier phases of the template signal, thereby explaining the observed structural similarity. Additionally, we prove that the convergence rate of Fourier phases is inversely proportional to the number of noise observations and, in the high-dimensional regime, to the Fourier magnitudes of the template signal. Moreover, in the high-dimensional regime, the EfN estimator converges to a scaled version of the template signal. This work not only deepens the theoretical understanding of the EfN phenomenon but also highlights potential pitfalls in template matching techniques and emphasizes the need for careful interpretation of noisy observations across disciplines in engineering, statistics, physics, and biology.
@@ 1 Introduction
While the EfN phenomenon has been analyzed in prior work (see Section 3 for more details), a comprehensive theoretical understanding of the EfN model remains limited. This work contributes to filling that gap by rigorously analyzing the relationship between the reconstructed signal and the underlying template. The term ’Einstein from Noise’ was popularized in [44], where the authors illustrated the phenomenon using an image of Einstein as the template signal. However, the underlying effect had been observed earlier in the cryogenic electron microscope literature (see, for instance, [46], and further details in Section 3). In this work, we refer to the average of the aligned pure noise observations as the EfN estimator. A detailed formulation of the problem is provided in Section 2.
Figure 1 illustrates the EfN process, which consists of two key stages. First, the observations are aligned with the template signal to achieve optimal alignment. Then, the aligned observations are averaged. The result is the EfN estimator, which shares a structural resemblance to the template image, though it is not an identical reproduction.
Figure 1: Einstein from Noise. The EfN estimator consists of three stages: (1) finding the index of the maximum of the cross-correlation ($\hat{\mathsf{R}}_{i}$) between the $i$-th noise signal ($n_{i}$) and the template signal (e.g., Einstein’s image); (2) cyclically shifting the noise signal by $\hat{\mathsf{R}}_{i}$; (3) averaging the shifted noise signals. In this paper, we characterize the relationship between the output of this process—the EfN estimator—and the template signal.
@@
The central results of this work are as follows. Our first result, stated in Theorem 4.1, shows that the Fourier phases of the EfN estimator converge to the Fourier phases of the template signal, as the number of noisy observations (denoted by $M$) converges to infinity. However, it is important to note that the EfN estimator’s Fourier magnitudes do not necessarily converge to those of the template signal. We also show that the Fourier phases’ mean squared error (MSE) decays to zero with a rate of $1/M$. Since the Fourier phases are responsible for the formation of geometrical image elements, such as contours and edges [36, 45], this clarifies why the resulting EfN estimator image exhibits a structural similarity to the template, but not necessarily a full recovery. Our second result, stated in Theorem 4.3, proves that in the high-dimensional regime, where the dimension of the signal diverges, the convergence rate of the Fourier phases is inversely proportional to the square of the Fourier magnitudes of the template signal. In this case, the Fourier magnitudes of the EfN estimator converge to a scaled version of the template’s Fourier magnitudes.
While Theorems 4.1 and 4.3 are proved under the assumption of white Gaussian noise, we also extend our analysis to more general noise models. In particular, we show that, although the convergence results in Theorems 4.1 and 4.3 do not necessarily hold under arbitrary noise statistics, several structural properties of the EfN estimator persist. First, in Proposition 5.1, we show that the EfN estimator remains positively correlated with the template for arbitrary noise statistics, even when the Fourier phases do not converge. Since the correlation between images often implies visual resemblance, this explains why the EfN estimator still exhibits structural similarity to the template. Second, in Theorem 5.2, we show that in the high-dimensional limit, if the noise signal is independent and identically distributed (i.i.d) (not necessarily Gaussian), then the same phase convergence behavior observed in the white Gaussian case still holds. Finally, in Proposition 5.4, we demonstrate that if the noise signal is Gaussian with circular symmetry, then the conclusions of Theorem 4.1 remain valid, even though the noise is not white.
@@
The remainder of this paper is organized as follows. Section 2 provides a detailed formulation of the problem. Section 3 discusses the connection between the EfN problem and single-particle cryo-electron microscopy (cryo-EM), the primary motivation for this work, and presents supporting empirical demonstrations. Our main theoretical results for white Gaussian noise observations, Theorems 4.1 and 4.3, are stated in Section 4. Extensions of these results to noise models beyond white Gaussian noise are presented in Section 5. Finally, we conclude with a discussion and outlook in Section 6.
2 Problem Formulation and Notation
Notations.
Throughout the rest of this paper, we use $\xrightarrow[]{{\cal D}}$, $\xrightarrow[]{{\cal P}}$, $\xrightarrow[]{\mathsf{a.s.}}$, and $\xrightarrow[]{{\cal L}^{p}}$, to denote the convergence of sequences of random variables in distribution, in probability, almost surely, and in ${\cal L}^{p}$ norm, respectively. Inner products in the Euclidean space between vectors $a$ and $b$ are written as either $a^{\top}b$ or $\langle a,b\rangle$.
Problem formulation.
$\text{(Postulated model)}\quad y_{i}=\mathcal{T}_{\ell_{i}}\cdot x+n_{i},$ (2.1)
where $\mathcal{T}_{\ell}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the cyclic shift operator defined by $\mathcal{T}_{\ell}z]_{r}\triangleq z_{(r-\ell)\bmod d}$ for all $z\in\mathbb{R}^{d}$ and indices $0\leq r\leq d-1$, and $n_{i}\sim\mathcal{N}(0,\sigma^{2}I_{d\times d})$ are i.i.d. Gaussian noise vectors. In reality, however, there is no underlying signal: the observations consist entirely of white Gaussian noise. That is, the true data-generating process follows the underlying model:
$\text{(Underlying model)}\quad y_{0},y_{1},\ldots,y_{M-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(\mathbf{0},\sigma^{2}I_{d\times d}),$ (2.2)
where $M$ denotes the number of observations. Since the data consists purely of white Gaussian noise, we will explicitly write $y_{i}=n_{i}$ to emphasize this fact.
$\hat{\mathsf{R}}_{i}\triangleq\underset{0\leq\ell<d}{\arg\max}\,\langle n_{i},\mathcal{T}_{\ell}x\rangle,$ (2.3)
where $n_{i}$ is the $i$-th noise observation, and $\mathsf{\hat{R}}_{i}$ defines the optimal cyclic shift that aligns the template signal $x$ with the noise observation $n_{i}$ in terms of cross-correlation.
$\hat{x}\triangleq\frac{1}{M}\sum_{i=0}^{M-1}{\mathcal{T}_{\mathsf{-\hat{R}}_{i}}}n_{i},$ (2.4)
where $\mathcal{T}_{\mathsf{-\hat{R}}_{i}}n_{i}$ represents the noise observation $n_{i}$ aligned by applying the inverse cyclic shift $\mathsf{\hat{R}}_{i}$ to best match the template signal. Throughout the text, we refer to $\hat{x}$ as the EfN estimator.
The EfN phenomenon states that, at least empirically, $\hat{x}$ and $x$ appear “close” in some sense; our goal is to understand this phenomenon mathematically. To that end, we will consider the two asymptotic regimes: the first corresponds to the classical setting where the number of observations $M\to\infty$ while the dimension $d$ is fixed (i.e., the number of observations diverges to infinity and the template vector dimension is fixed); the second is the high-dimensional regime, where $d\to\infty$ after $M\to\infty$. (i.e., both the number of observations and the dimension of the template signal diverge).
Fourier space notation.
where $j\triangleq\sqrt{-1}$, and $0\leq k\leq d-1$. Accordingly, we let $\mathsf{X}$, $\hat{\mathsf{X}}$, and $\mathsf{N}_{i}$, denote the DFTs of $x$, $\hat{x}$, and $n_{i}$, respectively, for $0\leq i\leq M-1$. These DFT sequences can be equivalently represented in the magnitude-phase domain as follows,
for $0\leq i\leq M-1$, where $\left|\mathsf{X}[k]\right|$, $\left|\mathsf{\hat{X}}[k]\right|$, and $\left|\mathsf{N}_{i}[k]\right|$ are the $k$-th Fourier component magnitudes of the template signal, the EfN estimator, and the $i$-th noise observation, respectively. Similarly, $\phi_{\mathsf{X}}[k]$, $\phi_{\mathsf{\hat{X}}}[k]$, and $\phi_{\mathsf{N}_{i}}[k]$ represent the corresponding $k$-th Fourier phases. Note that the random variables $\left\{|\mathsf{N}_{i}[k]|\right\}_{k=0}^{d/2}$ and $\left\{\phi_{\mathsf{N}_{i}}[k]\right\}_{k=0}^{d/2}$ are two independent sequences of i.i.d. random variables, such that, $|\mathsf{N}_{i}[k]|\sim\mathsf{Rayleigh}\left({\sigma^{2}}\right)$ has Rayleigh distribution, and the phase $\phi_{\mathsf{N}_{i}}[k]\sim\mathsf{Unif}[-\pi,\pi)$ is uniformly distributed over $\pi,\pi)$.
for $k=0,1,\ldots,d-1$, where $\left|\mathsf{N}_{i}[k]\right|$ and $\phi_{\mathsf{N}_{i}}[k]$ are defined in (2.6). It is important to note that the expression above will converge to zero without the last term that captures the dependency in $\mathsf{\hat{R}}_{i}$—the location of the maximum correlation. This term reflects the fundamental properties of the EfN process and its dependency on the template signal, as well as the connections between the different spectral components. We denote by $\mathbb{E}{\left|\phi_{\mathsf{\hat{X}}}[k]-\phi_{\mathsf{X}}[k]\right|^{2}}$ the MSE of the Fourier phases of the $k$-th spectral component.
Assumptions.
Throughout this paper, we assume that the template signal $x$ is normalized, i.e., $\left\|x\right\|_{2}^{2}=1$, where $\left\|\cdot\right\|_{2}$ is the Euclidean norm, and further assume that its Fourier transform in non-vanishing, except possibly at the DC (zero-frequency) component. The first assumption is used for convenience and does not alter (up to a normalization factor) our main results in Theorems 4.1 and 4.3. The second assumption is essential for the theoretical analysis of the EfN process and is expected to hold in many applications, including cryo-EM. A similar assumption is frequently taken in related work, e.g., [8, 38, 10]. It is worth noting that since the Fourier transform of $x$ is assumed to be non-vanishing, the maximizing shift $\hat{\mathsf{R}}_{i}$ in (2.3) is almost surely unique. In addition, without loss of generality, we assume that the signal length $d$ is even.
@@ 3 Cryo-EM and Empirical Demonstration
Cryo-EM is a powerful tool of modern structural biology, offering advanced methods to visualize complex biological macromolecules with ever-increasing precision. One of its central advantages lies in its capability to resolve the structures of proteins that are hard to crystallize in traditional methods, especially in a near-physiological environment, see e.g., [35, 48]. This advantage enables researchers to delve into the dynamic behaviors of proteins and their complexes, shedding light on fundamental biological processes.
Single-particle cryo-EM uses electron microscopy to reconstruct 3D structures from 2D tomographic projection images [9]. Typically, the 3D reconstruction involves two main steps: detecting and extracting single particle images using a particle picking algorithm, [42, 22, 11, 21], and then reconstructing the 3D density map [41, 40]. Most detection algorithms use template-matching techniques, which can introduce bias if improper templates are chosen, especially in low signal-to-noise ratio (SNR) conditions, which is the standard scenario in cryo-EM.
@@
As introduced in the previous section, the EfN phenomenon depends on several key parameters: (1) the number of observations, denoted by $M$; (2) the dimension of the signal, denoted as $d$ (for example, the number of pixels in Einstein’s image); and (3) the statistical properties of the template signal, and in particular its power spectral density (PSD). To demonstrate the dependency on these parameters and provide insight into our main results, Figures 2 and 3 show the convergence of the EfN estimator. Both figures were generated according to the procedure outlined in Section 2. When referring to Monte Carlo trials, it means that the EfN estimator procedure, as specified in (2.4), was executed multiple times (the number of Monte-Carlo trials), each trial with fresh data.
Figure 2 shows the EfN estimator as a function of $M$. Figure 2(a) shows that as $M$ increases, the EfN estimator becomes more structurally similar to the template Einstein image. Indeed, Figure 2(b) shows that as the number of observations $M$ grows, the MSE between the Fourier phases of the template image and the corresponding Fourier phases of EfN estimator decreases. Figure 2(c) highlights that the convergence rate is proportional to $1/M$, with a faster convergence rate for stronger spectral components.
Figure 3 illustrates the impact of the template signal’s PSD on the cross-correlation between the template signal and the EfN estimator. Notably, a flatter PSD (i.e., a faster decay of the auto-correlation) leads to a higher correlation between the template and the estimator signals. These empirical results are proved theoretically in Theorems 4.1 and 4.3.
@@
The EfN phenomenon extends to various applications employing template matching, whether through a feature-based or direct template-based approach. For instance, template matching holds significance in computational anatomy, where it aids in discovering unknown diffeomorphism to align a template image with a target image [15]. Other areas include medical imaging processing [1], manufacturing quality control [3], and navigation systems for mobile robots [28]. This pitfall may also arise in the feature-based approach, which relies on extracting image features like shapes, textures, and colors to match a target image by neural networks and deep-learning classifiers [59, 34, 53, 30].
Figure 2: The impact of the number of noise observations on the EfN estimator. The EfN estimator is defined in real space by (2.4) and in Fourier space by (2.7). (a) The structural similarity between the EfN estimator and the template image increases as a function of the number of noise observations ($M$). (b) The mean-square-error (MSE) between the Fourier phases of the template image $\mathsf{X}[k_{1},k_{2}]$ and the EfN estimator $\hat{\mathsf{X}}[k_{1},k_{2}]$ for $100\leq k_{1},k_{2}\leq 100$, where $k_{1},k_{2}$ are the indices of the 2D DFT. The colors in the left panel in (b) represent the power spectral density (PSD) of the Einstein image, while the colors in the four right panels represent the MSE between the Fourier phases of the Einstein image and the EfN estimator, for each spectral component, with a varying number of observations ($M=200,500,1500,5000$). An increase in the number of observations leads to a lower MSE of the Fourier phases between the EfN estimator and the template signal. A similar trend can be seen with respect to the strength of the spectral components, i.e., stronger spectral components lead to lower Fourier phases MSE. (c) The convergence rate of the MSE between the Fourier phases of the EfN estimator and the Fourier phases of the template signal as a function of the number of observations across different frequencies. The MSE decays as $1/M$. In addition, stronger spectral components lead to lower MSE. Figures (b) and (c) were generated through 200 Monte-Carlo trials of the EfN process defined in (2.4).
Figure 3: The influence of the power-spectral-density (PSD) of the template signal on the correlation between the template and the EfN estimator. (a) Three images of the letter A are shown, with an increasing zero-padding ratio. As the zero-padding ratio increases, the PSD flattens, and the cross-correlation ($C$) between the template and the EfN estimator increases. This higher cross-correlation is evident in both the image background and the colors of the letter A. (b) Flatter PSDs lead to EfN estimators whose Fourier magnitudes are closer to those of the template image. The EfN estimators in these experiments were generated using $M=10^{5}$ observations.
@@
The EfN phenomenon has been investigated in earlier studies. In particular, it was shown that the ratios between the expected values of the Fourier coefficients of the EfN estimator and those of the template are real-valued [60, Chapter 5]. In this work, we build upon and significantly extend these results. Specifically, we establish the convergence of the EfN estimator to a non-vanishing signal, derive its convergence rate, analyze its behavior in the high-dimensional regime, and generalize the analysis to encompass a broader class of noise models beyond white Gaussian noise.
@@
A closely related work is that of Wang et al. [55], who conducted a rigorous statistical analysis of model bias in a different but complementary setting. They analyze the effects of selectively averaging only samples that exhibit the highest cross-correlation with a fixed reference signal (e.g., Einstein’s image). This selection mechanism introduces a bias toward the reference, and their analysis reveals a phase transition in the resulting reconstruction, governed by the number of samples, the signal dimension, and the size of the selected subset. Notably, their results show that a structured image can emerge even when averaging purely noisy data. A related but statistically distinct selection-based mechanism is studied in [7] in the context of template matching for particle picking and extraction: candidate (pure-noise) observations are filtered by thresholding their cross-correlation with a bank of templates, and downstream averages inherit a strong template imprint.
In contrast, our work investigates the EfN estimator in the absence of any selection step: all pure-noise observations are first aligned to a fixed template and then averaged. This difference in mechanism leads to different statistical behavior. Whereas template-based selection can yield averages that closely resemble the templates up to a global scale factor, we show that the non-selective EfN procedure exhibits a more structured form of alignment-induced bias, manifested most prominently through phase locking: the Fourier phases of the estimator converge to those of the template, even though the resulting average is generally not identical to the template. Closely related alignment-induced artifacts and their connection to Fourier-phase behavior have also been discussed in the multireference alignment literature; see, e.g., [43].